OpenStudy (lgbasallote):

just checking.. A sample of nitrogen gas has a volume of 32.4 L at 20°C. The gas is heated to 220ºC at constant pressure. What is the final volume of nitrogen? A) 2.94 L B) 19.3 L C) 31.4 L D) 54.5 L E) 356 L i got B

sam (.sam.):

Use \[\frac{V_1}{T_1}=\frac{V_2}{T_2}\]

OpenStudy (lgbasallote):

\[\frac{32.4 \times 493.15}{293.15} = 54.50?\]

sam (.sam.):

yes

OpenStudy (lgbasallote):

hmm wonder what went wrong a while ago o.O

sam (.sam.):

Just make sure to convert the degree Celsius to kelvin